OpenStudy (anonymous):

How would I solve this? 6x=x+5

OpenStudy (anonymous):

6x=x+5 6x-x=5 x(6-1)=5 5x=x x=5/5 x=1

OpenStudy (solomonzelman):

Oh I se, my bad!

OpenStudy (anonymous):

I meant 5x=5 so x=5/5=1

OpenStudy (anonymous):

Thank you!